Suppose a firm has a short run total cost curve given by Tc=100+20Q+Q³. The corresponding short run marginal cost curve is MC=20+2Q. What is the equation for average variable cost? And What is the minimum level of average variable cost?


1
Expert's answer
2020-12-14T06:11:40-0500

TC = Q3 + 20Q + 100

TVC = Q3 + 20Q

TFC = 100

The equation for the average variable cost (AVC):

"AVC = \frac{TVC}{Q} \\

AVC = \frac{Q^3+20Q}{Q} = Q^2 + 20"

AVC reach its minimum if:

"AVC = MC \\

Q^2+20=20+2Q \\

Q = 2"

The minimum level of average variable cost Q = 2